OpenStudy (anonymous):

Solve by using tables. Give each answer to at most two decimal places. x^2 − 5x + 6 = 0 A, –5, 6 B. 2, –3 C.2, 3 D. –2, –3

OpenStudy (anonymous):

I think its b

OpenStudy (anonymous):

I think you mixed up a negative sign somewhere...

OpenStudy (anonymous):

x^2 - 5x + 6 = 0 Let's try x = 2: 2^2 - 5(2) + 6 = 4 - 10 + 6 = 0 <<<--- correct, so x = 2 is a solution. For x = -3: (-3)^2 - 5(-3) + 6 = 9 + 15 + 6 = 30... but obviously 30 is not equal 0, so x = -3 isn't a solution. Try x = 3 instead...